# Uniform irreducibility of Galois action on the $\ell$-primary part of Abelian $3$-folds of Picard type

### v1

@@ 1.5. Intertwining and an exponential sum
computing explicitly the right hand side as a rational function in $p^{-s}$, simultaneously justifying that $M_{s}$ does not have a pole at $s=0$.
The Lemma 1.11 implies the existence of a non-zero element $f_{y}\in\left(\operatorname{Ind}_{B}^{G}(\mu)\right)^{K_{T}}$ supported on $B\sigma_{y}K_{T}$, which we normalize by $f_{y}(\sigma_{y})=1$. Consider a $K^{\circ}$-flat section $f_{y,s}$ passing thru $f_{y,0}=f_{y}$, and computing $M_{s}(f_{y,s})$ for $\Re(s)\gg 0$.
@@
For all $y,y^{\prime}\in\mathbf{F}_{p}$, we have:
$M_{s}(f_{y^{\prime},s})(\sigma_{y})=\chi(-1)p^{-1}(1-p^{-1})(1-\mu_{s}(\gamma))^{-1}.$
Thus, to complete the proof, it suffices to show that $\pi_{n}^{I_{2}}\neq\{0\}$ or equivalently find $y\in\mathbf{F}_{p}^{\times}$ such that $f_{y}\notin\pi_{2}$.
Although we only solve this question in the case of a unitary group in $3$ variables, we feel that it deserves to be studied in greater generality for its own sake.
@@ 2. Galois representations for $3$-folds of Picard type
From this point onwards, we will use global notations. The local results of the previous section can be applied to the completion $E$ of $M$ at any prime number which does not split in that field.
We denote by $\mathbf{A}_{f}$ the ring of finite adeles of $\mathbf{Q}$, so that $\mathbf{A}=\mathbf{A}_{f}\times\mathbf{R}$.
2.1. Abelian $3$-folds of Picard type and Tate modules
@@
The action of $M$ splits the $3$-dimensional $k$-vector space $\operatorname{Lie}(A/k)$ in a direct sum of two sub-spaces: one on which the actions of $M$ and $k$ agree, and one on which they differ by the complex conjugation. We say that $A$ is of truly of Picard type if the pair of dimensions of these spaces, called the signature, equals $(2,1)$.
A polarization on $A$ is an isogeny $\theta:A\to A^{\vee}$, where $A^{\vee}$ denotes the dual abelian variety. By positivity, since $k$ is a field, the Rosati involution induced by $\theta$ on $\iota(\mathcal{O})$ is given by the complex conjugation (see [21, §21]). A polarization is called principal, if it is an isomorphism, and can can always be acquired over a finite extension of $k$.
@@
2.2. Shimura surfaces and families of abelian threefolds of Picard type
Let $(V,\langle\cdot,\cdot\rangle)$ be a $3$-dimensional (non-degenerate) hermitian space over $M$. The corresponding unitary similitude group $\widetilde{G}=\mathrm{GU}(V)$ is a reductive group over $\mathbf{Q}$ such that for any $\mathbf{Q}$-algebra $R$ one has:
@@ 2.2. Shimura surfaces and families of abelian threefolds of Picard type
where $\nu:\mathrm{GU}(V)\to\mathbf{G}_{m,\mathbf{Q}}$ is a homomorphism whose kernel is the unitary group $G=\mathrm{U}(V)$.
Note that any hermitian form in $3$ variables over a non-archimedean local field is isotropic, hence $\widetilde{G}(\mathbf{Q}_{p})$ is unique up to isomorphism, while at infinity $\langle\cdot,\cdot\rangle$ is uniquely determined by its signature, hence there are only two possibilities for $\widetilde{G}(\mathbf{R})$ (as opposite signatures define isomorphic groups). Hasse’s Principle applied to the semi-simple simply connected derived group $\widetilde{G}^{1}=\mathrm{SU}(V)$ implies then that, up to an isomorphism, there exists a unique quasi-split unitary group, denoted $\mathrm{GU}_{2,1}$, and a unique definite unitary group, denoted $\mathrm{GU}_{3,0}$.
We will now define the Shimura variety for the unitary similitude group $\widetilde{G}=\mathrm{GU}_{2,1}$ represented by the matrix $\left(\begin{smallmatrix}&&1\\ &\sqrt{-D}&\\ -1&&\end{smallmatrix}\right)$. The one for $\mathrm{GU}_{1,2}$ is its complex conjugate. The homomorphism of $\mathbf{R}$-algebraic groups
@@
satisfy the Shimura datum axioms for $\widetilde{G}$, hence for any open compact subgroups $\widetilde{K}$ of $\widetilde{G}(\mathbf{A}_{f})$ one can consider the Shimura surface
@@
where $\mathcal{H}\simeq\widetilde{G}(\mathbf{R})/\widetilde{K}_{\infty}$ is identified with the $\widetilde{G}(\mathbf{R})$-conjugacy classes of $\tilde{h}$. By a fundamental result of Shimura $Y_{\widetilde{K}}$ admits a canonical model over the reflex field $M$. As we will see, the connected components of $Y_{\widetilde{K}}$ are Picard modular surfaces, justifying the terminology.
For $\widetilde{G}$ anisotropic, can analogously define Shimura sets which are finite and therefore will not alter the uniformity of our results in §4.
@@
Henceforth, we will only consider abelian $3$-folds which are principally polarized and admit multiplication by $\mathcal{O}_{M}$, and we will refer to them simply as being of Picard type.
It must be noted that, even though each point of $Y_{\widetilde{K}}(\mathbf{C})$ is associated to an abelian $3$-fold of Picard type, there does not exist such a family over the entire $Y_{\widetilde{K}}(\mathbf{C})$ unless there is no point with extra automorphisms, in which case $Y_{\widetilde{K}}(\mathbf{C})$ would be a fine moduli space. In our cases of interest $\widetilde{K}$ is not neat, and therefore $S_{\widetilde{K}}$] is not a fine moduli space.
However, given an open compact subgroup $\widetilde{K}$, we claim that there is an abelian family of Picard type $A$ over any open subset $U$ of $Y_{\widetilde{K}}$ which contains no point with a non-trivial stabilizer. By [17, §2.3.4], the moduli stack $S_{\widetilde{K}}$ associated to this problem is an algebraic stack (for the étale topology), locally of finite type over the base which we may take to $\operatorname{Spec}(M)$. Moreover, by [17, §A.7.5], there is a canonical surjective morphism $\phi$ from $S_{\widetilde{K}}$ to the associated coarse moduli space $S_{\widetilde{K}}$], which in our notations is $Y_{\widetilde{K}}$. By [17, §7], $S_{\widetilde{K}}$] is an algebraic space and even a quasi-projective scheme. Moreover, by a general property of moduli stacks (see [22, Chap. 7]), $\phi$ is an isomorphism over the locus $U$ where there is no non-trivial automorphism, by which we mean it has no infinitesimal automorphism; analytically, this corresponds to points of $Y_{\widetilde{K}}(\mathbf{C})$ having no non-trivial stabilizers. Now $U$ is a priori an open subscheme of $S_{\widetilde{K}}$], but since it is where $\phi$ is an isomorphism, we get a canonical open $j:U\to S_{\widetilde{K}}$. This map tautologically yields the desired family $f:A\to U$ of abelian varieties of Picard type, whose existence is essential for our proof of the main results.
In analogy with Gross’ index $2$ subgroups of maximal compacts $K^{\circ}_{p}$ at ramifies primes introduced in §1.3, we consider the open compact subgroup
@@
Let $(A,\iota,\theta)$ be a principally polarized abelian $3$-fold of Picard type over $k$. For $v$ the prime of $M$ above $p\mid D$, the action of the absolute Galois group $\operatorname{Gal}_{k}$ on $A[v]$ factors through $\mathrm{GO}(3,\mathbf{F}_{p})$. Using the exceptional isomorphism $\mathrm{PGO}(3,\mathbf{F}_{p})\xrightarrow{\sim}\mathrm{SO}(3,\mathbf{F}_{p})\xrightarrow{\sim}\operatorname{PGL}_{2}(\mathbf{F}_{p})$, one defines its projectivization
(8) $\widetilde{\rho}_{A,p}:\operatorname{Gal}_{k}\to\operatorname{PGL}_{2}(\mathbf{F}_{p}).$

### v2

@@ 1.5. Intertwinings and exponential sums
as a rational function in $p^{-s}$, simultaneously justifying the absence of pole at $s=0$.
For any $y\in\mathbf{F}_{p}$, Lemma 1.10 yields a unique function $f_{y}\in\left(\operatorname{Ind}_{B}^{G}(\mu_{\lambda})\right)^{K_{T}}$ supported on $B\sigma_{y}K_{T}$ and normalized by $f_{y}(\sigma_{y})=1$. In addition, when $\chi$ is quadratic, we let $f_{\infty}$ be the unique function supported on $B\sigma_{\infty}K_{T}$ and normalized by $f_{\infty}(\sigma_{\infty})=1$. Consider a $K^{\circ}$-flat section $f_{y,s}$ passing through $f_{y,0}=f_{y}$.
Proposition 1.12.
For all $y,y^{\prime}\in\mathbf{F}_{p}\cup\{\infty\}$ and $\Re(s)\gg 0$, we have
$M_{s}(f_{y^{\prime},s})(\sigma_{y})=\chi(-1)\frac{(1-p)p^{-3-2s}}{1-p^{-1-2s}}\delta_{y,y^{\prime}}+p^{-3}\sum_{[z,x]\in N(\mathbf{F}_{p})}f_{y^{\prime},s}\left(\mathsf{w}[z,x]\sigma_{y}\right).$
Proof.
Recall that $z,x]=\left(\begin{smallmatrix}1&-\bar{z}&u\\ &1&z\\ &&1\end{smallmatrix}\right)$, where $u=\xi\cdot x-\tfrac{z\bar{z}}{2}$, so that $u+\bar{u}+z\bar{z}=0$. We first compute the intertwining integral over
[…]
a key observation being that $z,x]\in N(\mathbf{Z}_{p})$ if and only if $u\in\mathcal{O}$. It follows that for $z,x]\in N(\mathbf{Q}_{p}){\smallsetminus}N(\mathbf{Z}_{p})$ both $1/u$ and $\bar{z}/u$ belong to $\mathcal{P}$, yielding an explicit Iwasawa decomposition:
$\mathsf{w}n=\mathsf{w}[z,x]=\begin{pmatrix}&&1\\ &1&z\\  1&-\bar{z} &u\end{pmatrix}=\begin{pmatrix}\bar{u}^{-1}&\bar{z}/u&1\\ &-\bar{u}/u&z\\  &&u\end{pmatrix}\begin{pmatrix}1&&\\ z/\bar{u}&1&\\  1/u&-\bar{z}/u &1\end{pmatrix}$.
As the last term belongs to the principal congruence subgroup $K_{1}$ which is normalized by $\sigma_{y}\in K^{\circ}$ and contained in $K_{T}$, we deduce that $f_{y^{\prime},s}(\mathsf{w}n\sigma_{y})=\begin{cases}\mu_{\lambda,s}(\bar{u}^{-1},-\bar{u}/u,u)&\text{ if } y^{\prime}=y,\\ 0&\text{ if } y^{\prime}\neq y.\end{cases}$ Hence for $u\in p^{-m}\mathcal{O}^{\times}$, we have $f_{y^{\prime},s}(\mathsf{w}n\sigma_{y})=\mu_{\lambda,s}(\gamma)^{-m}\chi^{3}_{E}(\bar{u}^{-1})=(-1)^{m}\cdot p^{-(3+2s)m}\chi^{3}_{E}(u)$.
For $m$ odd, noticing further that $\chi_{E}(u)=\chi_{E}(\xi)=\chi(-1)$, we find
$\int\limits_{[p^{(1-m)/2}\mathcal{O},p^{-m}\mathbf{Z}_{p}^{\times}]}f_{y,s}(\mathsf{w}n\sigma_{y})dn=-\chi(-1)\cdot p^{-(3+2s)m}p^{m-1}(p^{m}-p^{m-1}),\text{ hence}$
(7) $\sum\limits_{m\geqslant 1\text{ odd}} \int\limits_{[p^{(1-m)/2}\mathcal{O},p^{-m}\mathbf{Z}_{p}^{\times}]}f_{y,s}(\mathsf{w}n\sigma_{y})dn=\frac{-\chi(-1)(p^{-2-2s}-p^{-3-2s})}{1-p^{-2-4s}}.$
We similarly find
(8) $\sum\limits_{m\geqslant 2\text{ even}}\int\limits_{[p^{(2-m)/2}\mathcal{O},p^{-m}\mathbf{Z}_{p}^{\times}]}f_{y,s}(\mathsf{w}n\sigma_{y})dn=\frac{\chi(-1)(p^{-4-4s}-p^{-5-4s})}{1-p^{-2-4s}}.$
Finally, for $m\geqslant 2$ even we compute
$\int\limits_{[p^{-\frac{m}{2}}\mathcal{O}^{\times},p^{-m}\mathbf{Z}_{p}]}f_{y,s}(\mathsf{w}n\sigma_{y})dn=p^{-(3+2s)m}\cdot p^{m-2}p^{m-1} \sum\limits_{[z_{0},x_{0}]\in\mathbf{F}_{p^{2}}^{\times}\times\mathbf{F}_{p}}\chi^{3}_{E}(z_{0}\bar{z}_{0}-2\xi x_{0}).$
As $\chi^{3}\neq\mathbf{1}$, $\sum\limits_{(z_{0},x_{0})\in\mathbf{F}_{p^{2}}^{\times}\times\mathbf{F}_{p}}\chi^{3}\left(\frac{z_{0}\bar{z}_{0}-2\xi x_{0}}{z_{0}\bar{z}_{0}+2\xi x_{0}}\right)=(p^{2}-1)\sum\limits_{x_{0}\in\mathbf{F}_{p}}\chi^{3}\left(\frac{1-\xi x_{0}}{1+\xi x_{0}}\right)=-\chi(-1)(p^{2}-1)$ yields
(9) $\sum\limits_{m\geqslant 2\text{ even}}\int\limits_{[p^{-\frac{m}{2}}\mathcal{O}^{\times},p^{-m}\mathbf{Z}_{p}]}f_{y^{\prime},s}(\mathsf{w}n\sigma_{y})dn=\frac{-\chi(-1)(p^{-3-4s}-p^{-5-4s})}{1-p^{-2-4s}}.$
Putting (7), (8) and (9) together, we find $\int\limits_{N(\mathbf{Q}_{p}){\smallsetminus}N(\mathbf{Z}_{p})}f_{y,s}(\mathsf{w}n\sigma_{y})dn=-\chi(-1)\frac{(p^{-2-2s}-p^{-3-2s})}{1-p^{-1-2s}}$.
Finally, as $f_{y^{\prime},s}$ is invariant by $K_{T}\supset K_{1}\supset N(p\mathbf{Z}_{p})$, we have
$\int\limits_{N(\mathbf{Z}_{p})}f_{y^{\prime},s}(\mathsf{w}n\sigma_{y})dn=p^{-3}\sum_{[z,x]\in N(\mathbf{Z}_{p})/N(p\mathbf{Z}_{p})}f_{y^{\prime},s}\left(\mathsf{w}[z,x]\sigma_{y}\right).\qed$
The coefficients of the intertwining automorphism of $\operatorname{Ind}_{\overline{B}}^{\overline{G}}(\overline{\mu}_{\lambda})^{\overline{T}}$ in the basis $(f_{y})_{y}$ are given by
$M_{y,y^{\prime}}=\sum_{[z,x]\in\mathbf{F}_{p^{2}}\times\mathbf{F}_{p}}f_{y^{\prime},0}\left(\mathsf{w}[z,x]\sigma_{y}\right).$
Corollary 1.13.
Letting $\mathscr{M}_{0}$ (resp. $\mathscr{M}_{0}^{\mathsf{w}}$) denote the matrices of $M_{0}$ (resp. $M_{0}^{\mathsf{w}}$) in the bases $(f_{y})_{y}$ and $(f^{\mathsf{w}}_{y})_{y}$, we have
$\mathscr{M}:=(M_{y,y^{\prime}})_{(y,y^{\prime})}=\chi(-1)p\cdot I_{p}+p^{3}\cdot\mathscr{M}_{0}=-\chi(-1)p^{2}\cdot I_{p}+p^{3}\cdot\mathscr{M}_{0}^{\mathsf{w}}.$
Proof.
The first equality is obtained by letting $s$ go to 0 in Proposition 1.12, while the second one is obtained by applying the same process to $\mu_{\lambda}^{\mathsf{w}}$, noting that $\mu_{\lambda,s}^{\mathsf{w}}=\mu_{\lambda,s-1}$, in particular $\overline{\mu}_{\lambda}=\overline{\mu}_{\lambda}^{\mathsf{w}}$. ∎
@@
For all $y\in\mathbf{F}_{p}$, one has $\sum\limits_{y^{\prime}\in\mathbf{F}_{p}}M_{y,y^{\prime}}=\sum\limits_{y\in\mathbf{F}_{p}}M_{y,y^{\prime}}=\chi(-1)p$. More importantly
(10) $\operatorname{Tr}(\mathscr{M})=\begin{cases}-\chi(-1)p&\text{ if } \chi^{2}\neq\mathbf{1}\quad(\text{and} \chi^{3}\neq\mathbf{1}),\\ 0&\text{ if } \chi\text{ quadratic}.\end{cases}$
Proof.
As $f_{y^{\prime}}(\mathsf{w}\sigma_{y})=0$, we can assume that $u\neq 0$. Letting $v=1/u$ and $w=1-\bar{z}/u$, we have
$\mathsf{w}[z,x]\sigma_{y}=\begin{pmatrix}-\tfrac{1}{2}+\xi y&1&1\\ -1-\tfrac{z}{2}+\xi yz&1+z&z\\  1+\bar{z}-\tfrac{u}{2}+\xi yu&u-\bar{z}&u\end{pmatrix}=\begin{pmatrix}\bar{u}^{-1}&*&*\\ &-\bar{u}/u&*\\  &&u\end{pmatrix}\begin{pmatrix}1&&\\ -\bar{w}&1&\\  v-w+\tfrac{1}{2}+\xi y&w&1\end{pmatrix}$
Hence $f_{y^{\prime}}(\mathsf{w}[z,x]\sigma_{y})\neq 0$ if and only if $v-w+\tfrac{1}{2}+\xi y=w\bar{w}(-\tfrac{1}{2}+\xi y^{\prime})\neq 0$, i.e.,
[…]
in which case it equals $\chi_{E}(u^{3}w)=\chi_{E}(w)\cdot\chi^{3}_{E}(\bar{v})$. Noting $(w-1)(\bar{w}-1)+v+\bar{v}=0$, so that for a fixed $w\neq 1$, the ratio $\tfrac{\bar{v}}{v}=\tfrac{u}{\bar{u}}=\frac{\xi x-(w-1)(\bar{w}-1)/2}{-\xi x-(w-1)(\bar{w}-1)/2}$ describes $\mathbf{F}^{1}_{p^{2}}{\smallsetminus}\{-1\}$, we find
$\sum_{y\in\mathbf{F}_{p}}M_{y,y^{\prime}}=\sum_{y^{\prime}\in\mathbf{F}_{p}}M_{y,y^{\prime}}=$
$=\sum_{w\neq 1} \chi_{E}(w)\sum_{x\in\mathbf{F}_{p}} \chi^{3}_{E}\left(\xi x-(w-1)(\bar{w}-1)/2\right)+\sum_{x\in\mathbf{F}_{p}^{\times}} \chi^{3}(-1)$
$=\sum_{w\neq 1} \chi_{E}(w)(-\chi(-1))+(p-1)\chi(-1)=p\chi(-1).$
To compute $\operatorname{Tr}(\mathscr{M})$ we observe that when $w\notin\mathbf{F}^{1}_{p^{2}}$, (11) imposes no condition on $v$, as we are summing over all $y\in\mathbf{F}_{p}$, and that when $w\in\mathbf{F}^{1}_{p^{2}}$, it imposes no condition on $y$ while determines uniquely $v=\tfrac{w-\bar{w}}{2}-\tfrac{(w-1)(\bar{w}-1)}{2}=w-1$ . As $v\neq 0$ implies $w\neq 1$, we find
$\sum_{y\in\mathbf{F}_{p}}M_{y,y}=-\chi(-1)\sum_{w\in\mathbf{F}_{p^{2}}^{\times}{\smallsetminus}\mathbf{F}^{1}_{p^{2}}} \chi_{E}(w)+p\sum_{w\in\mathbf{F}^{1}_{p^{2}}{\smallsetminus}\{1\}} \chi_{E}(w(\bar{w}-1)^{3})$
$=\chi(-1)\sum_{w\in\mathbf{F}^{1}_{p^{2}}} \chi^{2}(w)+p\sum_{w\in\mathbf{F}^{1}_{p^{2}}{\smallsetminus}\{1\}}\chi(-\bar{w})=\begin{cases}-p\chi(-1)&\text{ if } \chi^{2}\neq\mathbf{1},\\ \chi(-1)&\text{ if } \chi\text{ quadratic}.\end{cases}$
Finally, for $\chi$ quadratic, letting $w=-\tfrac{\bar{z}}{u}$, we find $\mathsf{w}[z,x]\sigma_{\infty}=\left(\begin{smallmatrix}\bar{u}^{-1}&*&*\\ &-\bar{u}/u&*\\  &&u\end{smallmatrix}\right)\left(\begin{smallmatrix}1&&\\ -\bar{w}&1&\\  u^{-1}+\xi&w&1\end{smallmatrix}\right)$. This implies that $f_{\infty}(\mathsf{w}[z,x]\sigma_{\infty})=0$ unless $w=z=0$ in which case it equals $\chi_{E}(\xi y)$, with $y\bar{y}=1+\xi^{-2}x^{-1}\neq 0$. Then one finds $M_{\infty,\infty}=\sum\limits_{x\in\mathbf{F}_{p}^{\times}}f_{\infty}\left(\mathsf{w}[0,x]\sigma_{\infty}\right)=-\chi(-1)$. Combining with the second case of the above equation, one finds $\operatorname{Tr}(\mathscr{M})=M_{\infty,\infty}+\sum\limits_{y\in\mathbf{F}_{p}}M_{y,y}=0$. ∎
Proof of Theorem 1.9.
As $\overline{\mu}_{\lambda}=\overline{\mu}_{\lambda}^{\mathsf{w}}$, Frobenius reciprocity yields $\operatorname{End}_{\overline{G}}\left(\operatorname{Ind}_{\overline{B}}^{\overline{G}}(\overline{\mu}_{\lambda})\right)=\mathbf{C}^{2}$. Hence $\operatorname{Ind}_{\overline{B}}^{\overline{G}}(\overline{\mu}_{\lambda})$ is a direct sum of two irreducible distinct $\overline{G}$-representations, denoted $\overline{\pi}_{n}$ and $\overline{\pi}_{2}$, on each of which the intertwining endomorphism $\mathscr{M}$ acts scalarly by Schur’s Lemma.
Furthermore, as $\overline{\pi}_{n}^{\overline{T}}$ and $\overline{\pi}_{2}^{\overline{T}}$ are naturally identified with $\pi_{n}^{K_{T}}=\ker(\mathscr{M}_{0}^{\mathsf{w}})$ and $\pi_{2}^{K_{T}}=\ker(\mathscr{M}_{0})$ (see (2)), it follows from Corollary 1.13 that the above mentioned eigenvalues are respectively $\chi(-1)p^{2}$ and $\chi(-1)p\cdot I_{p}$. Formula (10) can thus be rewritten as
$p\cdot\dim\pi_{n}^{K_{T}}-\dim\pi_{2}^{K_{T}}=-\frac{\chi(-1)}{p}\operatorname{Tr}(\mathscr{M})=\begin{cases}1&\text{ if } \chi^{2}\neq\mathbf{1}\quad(\text{and} \chi^{3}\neq\mathbf{1}),\\ 0&\text{ if } \chi\text{ quadratic},\end{cases}$
which combined with (6) yields $\dim\pi_{n}^{K_{T}}=1$. ∎
@@ 2. Galois representations for $3$-folds of Picard type
From this point onwards, we will use global notations from the introduction. The local results of the previous section can be applied to the completion $E$ of $M$ at any prime number which does not split in that field. We denote by $\mathbf{A}_{f}$ the ring of finite adeles of $\mathbf{Q}$, so that $\mathbf{A}=\mathbf{A}_{f}\times\mathbf{R}$.
2.1. Abelian $3$-folds and Shimura surfaces of Picard type
@@
A polarization on $A$ is an isogeny $\theta:A\to A^{\vee}$, where $A^{\vee}$ denotes the dual abelian variety. By positivity, since $k$ is a field, the Rosati involution induced by $\theta$ on $\iota^{0}(M)$ is given by the complex conjugation (see [22, §21]). A polarization is called principal, if it is an isomorphism, and it can always be acquired over a finite extension of $k$.
By an abelian $3$-fold of Picard type over $k$ we will mean a principally polarized abelian variety over $k$ of dimension $3$ having multiplication by $\mathcal{O}_{M}$ defined over $k$.
The action of $M$ splits the $3$-dimensional $k$-vector space $\operatorname{Lie}(A/k)$ in a direct sum of two sub-spaces: one on which the actions of $M$ and $k$ agree, and one on which they differ by the complex conjugation, the pair of their dimension being called the signature.
@@
Let $(V,\langle\cdot,\cdot\rangle)$ be a $3$-dimensional (non-degenerate) hermitian space over $M$. The unitary similitude group $\widetilde{G}=\mathrm{GU}(V)$ is the reductive group over $\mathbf{Q}$ characterized by the property that
@@
for any $\mathbf{Q}$-algebra $R$, where $\nu:\mathrm{GU}(V)\to\mathbf{G}_{m,\mathbf{Q}}$ is a homomorphism whose kernel is the unitary group $G=\operatorname{U}(V)$. Since any hermitian form in $3$ variables over a non-archimedean local field is isotropic, the group $\widetilde{G}(\mathbf{Q}_{p})$ is unique up to isomorphism, while at infinity $\langle\cdot,\cdot\rangle$ is uniquely determined by its signature, hence there are two possibilities for $\widetilde{G}(\mathbf{R})$ (as opposite signatures define isomorphic groups). Hasse’s Principle applied to the semi-simple simply connected derived group $G^{1}=\mathrm{SU}(V)$ implies then that, up to an isomorphism, there exists a unique quasi-split unitary group, denoted $\widetilde{G}=\mathrm{GU}(2,1)$, and a unique definite unitary group, denoted $\mathrm{GU}(3)$.
We will now introduce the Shimura variety for $\widetilde{G}$ represented by the matrix $\left(\begin{smallmatrix}&&1\\ &\sqrt{-D}&\\ -1&&\end{smallmatrix}\right)$. The homomorphism
@@
of $\mathbf{R}$-algebraic groups satisfies Deligne’s axioms [8] for a Shimura variety, hence for any open compact subgroups $\widetilde{K}$ of $\widetilde{G}(\mathbf{A}_{f})$ one can consider the Shimura surface
@@
where $\mathcal{H}$ is identified with the $\widetilde{G}(\mathbf{R})$-conjugacy classes of $\tilde{h}$. By a fundamental result of Shimura, $Y_{\widetilde{K}}$ admits a canonical model over the reflex field $M$. The connected components of $Y_{\widetilde{K}}$ are Picard modular surfaces.
For the anisotropic form $\mathrm{GU}(3)$, one can analogously define Shimura sets which are finite and therefore will not alter the uniformity of our results in §3.
@@
Let $(A,\iota,\theta)$ be an abelian $3$-fold of Picard type over a number field $k$. The action of $\operatorname{Gal}_{k}$ on the adelic Tate module $V_{f}A$, together with the choice of $\eta$ as above, yields a continuous homomorphism:
(12) $\rho_{A,f}:\operatorname{Gal}_{k}\longrightarrow\widetilde{G}(\mathbf{A}_{f}).$
[…]
In analogy with the index $2$ Gross subgroups $\widetilde{K}^{\prime\prime}_{p}$ of maximal compacts $K^{\circ}_{p}$ at ramified primes introduced in §1.3, we consider the open compact subgroup
@@
Let $(A,\iota,\theta)$ be a principally polarized abelian $3$-fold of Picard type over $k$. For $v$ the prime of $M$ above $p\mid D$, the action of $\operatorname{Gal}_{k}$ on $A[v]$ factors through $\mathrm{GO}(3,\mathbf{F}_{p})$. Using the exceptional isomorphism $\mathrm{PGO}(3,\mathbf{F}_{p})\xrightarrow{\sim}\mathrm{SO}(3,\mathbf{F}_{p})\xrightarrow{\sim}\operatorname{PGL}(2,\mathbf{F}_{p})$, one defines its projectivization
(15) $\widetilde{\rho}_{A,p}:\operatorname{Gal}_{k}\to\operatorname{PGL}(2,\mathbf{F}_{p}).$